OpenStudy (anonymous):

Find the sum 7 divided by (x+3) + 9 divided by (x-3)

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bc}\] \[\frac{7}{x+3}+\frac{9}{x-3}=\frac{7(x-3)+9(x+3)}{(x+3)(x-3)}\]